Hey Siri to cycle start from the same point ride in opposite directions. One cyclist rides twice as fast as the other. In. 2 h, they are 54mi apart. Find the rate of each cyclist.
Answer by ikleyn(52775)

.

If x is the rate of the slower cyclist, then 2x is the rate of the other.


The total distance is the sum of partial distances


    2x + 2 *(2x) = 54.


    6x = 54


    x = 54/6 = 9 miles per hour.


The two rates are 9 mph and 18 mph.     ANSWER

Solved.

--------------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.